# Entanglement behavior and localization properties in monitored fermion systems

### v1

@@ Abstract
We study the asymptotic bipartite entanglement in various integrable and nonintegrable models of monitored fermions. We find that, for the integrable cases, the entanglement versus the system size is well fitted, over more than one order of magnitude, by a function interpolating between a linear and a power-law behavior. Up to the sizes we are able to reach, a logarithmic growth of the entanglement can be also captured by the same fit with a very small power-law exponent. We thus propose a characterization of the various entanglement phases using the fitting parameters. For the nonintegrable cases, as the staggered $t$-$V$ and the Sachdev-Ye-Kitaev (SYK) models, the numerics prevents us from spanning different orders of magnitude in the size, therefore we fit the asymptotic entanglement versus the measurement strength and then look at the scaling with the size of the fitting parameters. We find two different behaviors: for the SYK we observe a volume-law growth, while for the $t$-$V$ model some traces of an entanglement transition emerge. In the latter models, we study the localization properties in the Hilbert space through the inverse participation ratio, finding an anomalous delocalization with no relation with the entanglement properties. Finally, we show that our function also fits very well the fermionic logarithmic negativity of a quadratic model in ladder geometry, with stroboscopic projective measurements.
@@ I Introduction
Entanglement [1, 2] is a fundamental concept in quantum mechanics that has been widely exploited to characterize many-body quantum systems. It can spotlight the presence of quantum phase transitions [3] and the existence of topological boundary modes [4, 5, 6, 7], or can distinguish the unitary dynamics of thermalizing, integrable, and many-body localized quantum systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Moreover, in short-range systems undergoing a unitary evolution, it asymptotically attains a value that scales proportionally with the system size (volume law).
@@
Recently, attention has shifted to understanding the effects of an external monitoring on the asymptotic behavior of the entanglement. Intuitively, in the presence of projective measurements, the system collapses onto an eigenstate of the measurement operator and the entanglement remains constant with the system size (area law, in one-dimensional systems). However this scenario may change when undergoing both a measurement processes and a Hamiltonian evolution. Indeed, the interplay between the entangling effect of the Hamiltonian dynamics and the disentangling role of measurements leads to a variety of dynamical phases, characterized by peculiar entanglement behaviors. This gives rise to the entanglement transitions that have been identified in a variety of models, spanning from quantum circuits [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], to integrable or solvable [39, 40, 41, 42, 43, 44, 24, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] and nonintegrable [67, 68, 69, 70, 71, 72, 73, 74] Hamiltonian systems.
@@
Focusing on free-fermion systems in the presence of local weak measurements, a crossover from a phase in which the asymptotic bipartite entanglement entropy (EE) grows logarithmically with the system size to a phase in which it remains constant has been observed. The numerical evidence for a transition has been analytically confirmed in the case of a $\mathbb{Z}_{2}$ symmetry [48] and challenged for a $U(1)$ symmetry [49], suggesting that the logarithmic increase ceases for a size that is exponentially large in the inverse coupling. In this same context, when considering nonlocal weak measurements (e.g., power-law decaying measurement operators) or particular lattice geometries, one numerically observes transitions between three distinct situations: A volume-law, an intermediate subvolume-law, and an area-law entanglement phase [75, 76]. This is similar to what happens when the dynamics is only induced by random measurements of nonlocal strings [77, 78, 79].
@@
Here we look at this kind of transitions from a different unifying perspective, considering a variety of monitored fermionic systems (including the ones mentioned above) coupled to the environment through a quantum-state-diffusion monitoring process. We find that, in all cases, the asymptotic stationary value of the bipartite EE depends on the system size $L$ in a way that is remarkably well fitted, for all the sizes we study, by the function
@@
This behaves linearly at small $L$ and becomes a power law at large sizes, suggesting that the monitored system keeps being still in some way correlated, with a power-law increasing correlation length. Even though we lack a theoretical interpretation, we believe that this observation can contribute in the direction of a well stated description of the entanglement transitions.
The paper is divided in three parts. We first focus on integrable monitored fermionic models: (i) the tight-binding chain with onsite dephasing, (ii) the Kitaev chain with onsite dephasing, and (iii) the Kitaev chain with long-range dissipators. The EE has been already investigated in the literature, for these models. Here we go one step further and apply the fit in Eq. (1) to its dependence on the system size, finding a good agreement with the numerical data ranging from $L\sim 10$ to few hundreds. Moreover, from the value of the parameter $b$ obtained by the fit, we are able to identify the volume-law, subvolume-law, and area-law regimes that are believed to emerge in these models. For the range of system sizes we are able to access, the logarithm-law regime can be glimpsed by a power law with a very small exponent.
@@
In the second part, we extend our analysis to: (iv) the $t$-$V$ staggered model and (v) the Sachdev-Ye-Kitaev (SYK) model [80, 81], which have been recently considered in the context of entanglement transitions [82, 83]. Although their nonintegrability prevents us from accessing large sizes, we have worked out some scaling by fitting the asymptotic entanglement at fixed $L$ versus the coupling $\gamma$ to the environment. Then, we look at the scaling of the fitting parameter with the system sizes, finding that the $L$-dependence of Eq. (1) is recovered. Collecting all the results, we obtain that for the fully chaotic SYK model, the EE obeys a volume law at any measurement strength. Conversely, in the $t$-$V$ staggered model, we find traces of an entanglement transition in $\gamma$. We then address the localization properties through the time- and realization-averaged logarithm of the inverse participation ratio (IPR) in the Hilbert space. We find this quantity scales linearly with the logarithm of the dimension of the Hilbert space, with a slope that depends on $\gamma$. Its value predicts neither perfect delocalization nor perfect localization, but rather an anomalous delocalization, akin to multifractal behavior. This qualitative picture holds also when moving to the integrable limit and is independent of the entanglement behavior, suggesting that localization properties are not related to the entanglement transitions.
@@
In the third part, to witness the applicability of our procedure in a wider context, we show that it also applies to the fermionic logarithmic negativity (FLN). To do so, we focus on a ladder fermionic model undergoing projective measurements at discrete times [76, 84]. As for the EE, we find that the asymptotic FLN versus the system size is well described by Eq. (1), thus we are able to recognize the different dynamical regimes of the entanglement through the behavior of the FLN.
@@
The paper is organized as follows. In Sec. II we briefly recall the Lindblad description of monitored fermionic systems, focusing in particular on the quantum state diffusion protocol. In Sec. III we define the asymptotic bipartite EE and describe the proposed function to characterize its behavior. Then we present our results for integrable (Sec. IV) and for nonintegrable models (Sec. V). Finally, in Sec. VI we focus on the FLN in a ladder fermionic model. Our conclusions are drawn in Sec. VII.
@@ II Monitored fermionic systems
We consider systems of $L$ spinless fermions, described by Hamiltonians which can be generically cast as the sum of a quadratic and (possibly) a quartic term $\hat{H}=\hat{H}^{(2)}+\hat{H}^{(4)}$, where we define
[…]
@@
We are interested in describing the dynamics in the presence of weak measurements of some Hermitian operator $\hat{m}_{j}$. As is known [85, 86, 87], a single realization of the measurement sequence can be described by the stochastic evolution of a pure state |$\ket{\psi(t)}$ (namely, a quantum trajectory). On average, the system is described by a density matrix $\rho_{t}=\overline{\ket{\psi(t)}\bra{\psi(t)}}$ (the overline indicates ensemble averaging over many trajectories) obeying a Lindblad master equation
@@
Except for Sec. VI (whose details are given later), we implement a composite dynamics given by (i) an Hamiltonian evolution following a quantum quench and by (ii) a process of continuous measurement of the operator $\hat{m}_{j}$ (with $j=1,\dots,L$). The dynamics along each trajectory can be obtained by integrating the stochastic equation [85, 86, 87]
$d\ket{\psi(t)}$ $=$ $-\Big{[}i\hat{H}+\sum_{j}\frac{\gamma}{2}\big{(}\hat{m}_{j}-\braket{m_{j}}_{t}\big{)}^{2}\Big{]}dt\ket{\psi(t)}$ (6)
$+\Big{[}\sum_{j}\sqrt{\gamma}\,\big{(}\hat{m}_{j}-\braket{\hat{m}_{j}}_{t}\big{)}\,dW_{t}^{j}\Big{]}\ket{\psi(t)},$
where $\braket{\,\cdot\,}_{t}\equiv\braket{\psi(t)}{\cdot}{\psi(t)}$, while $W_{t}^{j}$ are independent Wiener processes (for $j=1,\ldots,L$). The state |$\ket{\psi(t)}$ along each trajectory appearing in Eq. (6) is called the unraveled state. We can discretize the evolution time with steps of length $\delta t$ and Trotterize the evolution. In what follows we consider measurement operators having the property
@@
|$\ket{\psi(t+\delta t)}\!\approx\mathcal{C}\ e^{\sum_{j}\!\!\big{[}\delta W^{j}_{t}+(2\braket{\hat{m}_{j}}_{t}-q_{j})\gamma\delta t\big{]}\hat{m}_{j}}e^{-i\hat{H}\delta t}\!\ket{\psi(t)}\!\,,$ (8)
@@
|$\ket{\psi(0)}=\prod_{j=1}^{L/2}\hat{c}_{2j}^{\dagger}\ket{\Omega},$ (9)
where |$\ket{\Omega}$ is the vacuum state for the $\hat{c}_{j}$-fermions. Note that, in general, this is not the ground state of the Hamiltonian $\hat{H}$ inducing the unitary part of the dynamics, so in this sense we are applying a quantum quench.
@@ III Asymptotic averaged bipartite entanglement entropy
being $\rho_{A}(t)=\mathrm{Tr}_{B}\big{[}\ket{\psi(t)}\bra{\psi(t)}\big{]}$ the reduced density matrix of subsystem $A$. Provided the global system is in a pure state [|$\ket{\psi(t)}$, in this case], the quantity $S_{\ell}(t)$ is a good measure of the entanglement between $A$ and $B$ and is usually referred to as the bipartite EE. Then, we average over many quantum trajectories
@@
Notice that this operation is different from evaluating the von Neumann entropy over the average state $\rho_{t}=\overline{\ket{\psi(t)}\bra{\psi(t)}}$ which, besides that, would also not be a proper measure of the entanglement. Finally, we fix $\ell$ to be a fixed fraction of $L$ (in particular we consider either $\ell=L/2$ or $\ell=L/4$) and estimate the asymptotic long-time value of $\overline{S_{\ell}(t)}$ by performing a suitable time average:
@@
Here $[t_{0},t_{f}]$ is an appropriate time window in which the behavior of $\overline{S_{\ell}(t)}$ has attained stationary. We perform the average over quantum trajectories numerically, over a finite number of realizations that we fix as $N_{\rm r}=48$. The error bars for our data are evaluated as the standard error (root-mean square deviation divided by $\sqrt{N_{\rm r}}$).
@@
[…]
In what follows, we fit $\overline{S}_{\ell}$ versus $L$ for different models with Eq. (1). The most interesting result is that this function seems to fit our numerics very well, independently of the considered model. In particular, in the next section we specialize to integrable models, while in Sec. V we focus on nonintegrable models. For the latter, due to the small attainable system sizes ($L\lesssim 20$), it is more convenient to fit $\overline{S}_{\ell}$ versus the coupling $\gamma$ with the environment. We choose a generalized Lorentzian function and, in the end, we find that the parameters of the fit scale with the system size, in such a way that the form in Eq. (1) is recovered (details are provided in Sec. V.3). Finally, in Sec. VI, we switch to a model with a more complicated geometry and a stroboscopic evolution, for which we consider the entanglement between two portions of a part of the whole system: as in that case the relevant part of the system is described by a mixed state, we resort to a proper entanglement monotone such as the FLN. To keep the presentation more accessible, we postpone all the required definitions to that section.

### v2

@@ Abstract
We study the stationary bipartite entanglement in various integrable and nonintegrable models of monitored fermions evolving along quantum trajectories. We find that, for the integrable cases, the entanglement versus the system size is well fitted, over more than one order of magnitude, by a function interpolating between a linear and a power-law behavior. Up to the sizes we are able to reach, a logarithmic growth of the entanglement can also be captured by the same fit with a very small power-law exponent. For the nonintegrable cases, such as the staggered $t$-$V$ and the Sachdev-Ye-Kitaev (SYK) models, the numerical limitations prevent us from spanning different orders of magnitude in the system size. Here we fit the asymptotic entanglement versus the measurement strength with a generalized Lorentzian, finding a very good agreement, and then look at the scaling with the size of the fitting parameters. We find two different behaviors: for the SYK we observe a linear increase with the system size, while for the $t$-$V$ model we see the emergence of traces of an entanglement crossover. In the latter models, we study the localization properties in the Hilbert space through the inverse participation ratio, finding an anomalous-delocalization behavior with no relation with the entanglement properties. Finally, we show that our function also fits very well the system-size dependence of the fermionic logarithmic negativity of a quadratic model in a two-leg ladder geometry, with stroboscopic projective measurements.
@@ I Introduction
Entanglement [1, 2] is a fundamental concept in quantum mechanics that has been widely exploited to characterize many-body quantum systems. It can spotlight the presence of quantum phase transitions [3] and the existence of topological boundary modes [4, 5, 6, 7], or can distinguish the unitary dynamics of thermalizing, integrable, and many-body localized quantum systems [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Moreover, in short-range systems undergoing a unitary evolution, it attains at long times a value that scales proportionally with the system size (volume law).
@@
Recently, attention has shifted to understanding the effects of an external monitoring on the long-time steady-state behavior of the entanglement. Intuitively, in the presence of projective measurements, the system collapses onto an eigenstate of the measurement operator and the entanglement remains constant with the system size (area law, in one-dimensional systems). However this scenario may change when undergoing both a measurement processes and a Hamiltonian evolution. Indeed, the interplay between the entangling effect of the Hamiltonian dynamics and the disentangling role of measurements leads to a variety of dynamical phases, characterized by peculiar entanglement behaviors. This gives rise to the entanglement transitions that have been identified in a variety of models, spanning from quantum circuits [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], to integrable or solvable [40, 41, 42, 43, 44, 45, 46, 25, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] and nonintegrable [69, 70, 71, 72, 73, 74, 75, 76] Hamiltonian systems.
@@
Focusing on free-fermion systems in the presence of local weak measurements, a crossover from a phase in which the steady-state bipartite entanglement entropy (EE) grows logarithmically with the system size to a phase in which it remains constant has been observed. The numerical evidence for a transition has been analytically confirmed in the case of a $\mathbb{Z}_{2}$ symmetry [50] and challenged for a $U(1)$ symmetry [51], suggesting that the logarithmic increase ceases for a size that is exponentially large in the inverse coupling, thus being, in the latter case, just a finite-size crossover. In the same context, when considering nonlocal weak measurements (e.g., power-law decaying measurement operators) or particular lattice geometries, one numerically observes transitions between three distinct situations: A volume-law, an intermediate subvolume-law, and an area-law entanglement phase [77, 78]. This is similar to what happens when the dynamics is only induced by random measurements of nonlocal strings [79, 80, 81].
@@
Here we look at these phenomena from a different perspective. We start considering some integrable monitored fermionic systems (including the ones mentioned above) coupled to the environment through a quantum-state-diffusion monitoring process. More specifically, we study (i) the tight-binding chain with onsite dephasing, (ii) the Kitaev chain with onsite dephasing, and (iii) the Kitaev chain with long-range dissipators. For these models the EE has been already investigated in the literature. Here we consider the problem from a slightly different perspective and find that the steady-state value of the bipartite EE depends on the system size $L$ in a way that is well fitted, for all the sizes we study, by the function
@@
The rationale behind this fitting function is the following. For small system sizes, one expects a linear increase of the entanglement entropy with $L$. The reason is that, in the absence of dissipation, the entanglement entropy increases linearly in time, due to the local-Hamiltonian ballistic propagation of correlations [9, 82], and this propagation is stopped by the environment measurements, on average after a characteristic time scale. This is especially true in the integrable systems we focus on, due to the ballistic propagation of quasiparticles [83]. In the presence of dissipation, this propagation stops on average after a characteristic time scale, due to the environment measurements [46]. So, we expect to see a linear increase of the steady-state entanglement entropy with the size, if $L$ is smaller than $v_{0}t^{*}$, where $t^{*}$ is the characteristic time after which the propagation stops and $v_{0}$ is the quasiparticle propagation velocity. In fact, for $L<v_{0}t^{*}$, quasiparticles (and correlations they bring with them) have the time to saturate the full chain before their propagation is arrested, thus the long-time entanglement entropy is linear in $L$. In agreement with that, a linear increase in $L$ for small system sizes is exactly provided by Eq. (1) in the limit $L\ll C^{-1/b}$. On the opposite hand, for large system sizes ($L\gg C^{-1/b}$), Eq. (1) gives rise to a power-law behavior $f(L)\sim\frac{A}{C}L^{1-b}$. We should stress that the fit is reliable as long as the numerical points encompass a range with a maximum $L_{\rm max}$ larger or equal than the length scale
$L_{0}\equiv C^{-1/b}$ (2)
separating the short-range linear from the long-range power-law behavior. For all the fits we perform with Eq. (1), we check that this condition is verified. In the long-range power-law regime we find different behaviors (depending on the value of $b$) for different models.
In particular, in the tight-binding chain with onsite dephasing it is well known that the system asymptotically in the system size always obeys an area-law behavior [46, 51]. In this case therefore we fit the steady-state EE with the function Eq. (1) fixing $b=1$ so that the large-system-size behavior is an area law. We find a very good agreement between the fitting function and the numerical point, fully confirming the theoretical predictions. We numerically find that $L_{0}$ scales as a power law in $\gamma$, not so far from the $\gamma^{-1}$ predictions of Ref. [46].
The Kitaev chain with and long-range dissipators is peculiar as well. Here we previously found [77] a volume-law regime, an area-law regime, and an intermediate subvolume regime that impossible to fit logarithmically. We find that the fit with Eq. (1) works fine in the three regimes, providing an asymptotic linear increase ($b=0$) in the volume law regime, an asymptotic finite value ($b=1$) in the area law regime, and an asymptotic power-law behavior ($0<b<1$) in the intermediate regime. In the intermediate regime $b$ (and the power-law exponent 1b1-b) displays a plateau. A similar power-law behavior, called “superdiffusive behavior”, has already been observed in a model [84] where symmetries allow quasiparticles to propagate long enough to enhance the correlations and increase the steady-state entanglement. In our model correlations are instead physically created by the long-range measured operators decaying as a power law. Moreover, we find that $L_{0}$ diverges when $\alpha$ tends to 1 from above. Indeed, for $\alpha<1$, the system displays a volume-law behavior of entanglement and quite consistently the scale $L_{0}$ separating volume from power-law behavior tends to infinity.
We point out that, at the present stage, our findings are only based on a purely numerical analysis, therefore Eq. (1) should not be exploited to classify the various entanglement phases. Despite this, as we shall see below, the remarkably good agreement of such fitting formula with the numerical data (even in the small-size range, differently from other proposed scaling Ansätze) suggests that Eq. (1) may motivate future investigations on the entanglement dynamics in monitored systems.
@@
Then we move on and extend our analysis to two nonintegrable models: (iv) the $t$-$V$ staggered model and (v) the Sachdev-Ye-Kitaev (SYK) model [85, 86], which have been recently considered in the context of entanglement transitions [87, 88]. Although their nonintegrability prevents us from accessing large sizes, we have worked out some scaling by fitting with a generalized Lorentzian the asymptotic entanglement at fixed $L$ versus the coupling $\gamma$ to the environment. Then, we look at the scaling of the fitting parameter with the system sizes, finding that the $L$-dependence of Eq. (1) is recovered. Collecting all the results, we obtain that for the fully chaotic SYK model, the EE linearly increases with the system size at any measurement strength. Conversely, in the $t$-$V$ staggered model, we find hints of an entanglement crossover in $\gamma$.
We also address the localization properties of these models. Analyses of localization/delocalization properties in some integrable monitored systems already exist [89, 67, 43]; here we extend this analysis to nonintegrable cases. Namely, we study the time- and realization-averaged logarithm of the inverse participation ratio (IPR) in the Hilbert space and find that this quantity scales linearly with the logarithm of the dimension of the Hilbert space, with a slope that depends on $\gamma$. Its value corresponds neither to perfect delocalization nor to perfect localization, but rather to an anomalous delocalization, akin to a multifractal behavior. This qualitative picture holds also when moving to the integrable limit and is independent of the entanglement behavior, suggesting that localization properties are not related to the entanglement transitions.
@@
To witness the applicability of our procedure in a wider context, in the third part we show that it also applies to the fermionic logarithmic negativity (FLN). To do so, we focus on a noninteracting fermionic model on a two-leg ladder, undergoing projective measurements at discrete times [78, 90]. As for the EE, we find that the asymptotic FLN versus the system size is well described by Eq. (1), thus we are able to recognize traces of the different dynamical regimes of the entanglement through the behavior of the FLN.
@@
The paper is organized as follows. In Sec. II we briefly recall the Lindblad description of monitored fermionic systems, focusing in particular on the quantum state diffusion protocol. In Sec. III we define the asymptotic bipartite EE and describe the proposed function to characterize its behavior. Then we present our results for integrable (Sec. IV) and for nonintegrable models (Sec. V). Finally, in Sec. VI we focus on the FLN in a ladder fermionic model. Our conclusions are drawn in Sec. VII. In the Appendix we provide further details on the stability of our fit (App. A), on the time traces of the trajectory-averaged EE (App. B), and on how to compute the FLN for the monitored noninteracting fermionic ladder (App. C).
@@ II Monitored fermionic systems
We consider systems of spinless fermions on a lattice with $L$ sites, described by Hamiltonians which can be generically cast as the sum of a quadratic and (possibly) a quartic term $\hat{H}=\hat{H}^{(2)}+\hat{H}^{(4)}$, where we define
$\hat{H}^{(2)}=$ $\sum_{i,j=1}^{L}\big(D_{ij}\,\hat{c}_{i}^{\dagger}\hat{c}_{j}+O_{ij}\,\hat{c}^{\dagger}_{i}\hat{c}^{\dagger}_{j}+\text{h.c.}\big)\,$, (3a)
$\hat{H}^{(4)}=$ $\sum_{i,j,k,l=1}^{L}\big(J_{ij,kl}\,\hat{c}_{i}^{\dagger}\hat{c}^{\dagger}_{j}\hat{c}_{k}\hat{c}_{l}+\text{h.c.}\big)\,.$ (3b)
@@
[…]
@@
Except for Sec. VI (whose details are given later), the results in Sec. IV and in Sec. V are obtained by implementing a composite dynamics given by (i) an Hamiltonian evolution following a quantum quench and by (ii) a process of continuous measurement of the operator $\hat{m}_{j}$ (with $j=1,\dots,L$). The dynamics, known as quantum state diffusion, along each trajectory can be obtained by integrating the stochastic equation [91, 92, 93]
[…]
@@
[…]
@@
[…]
@@ III Asymptotic averaged bipartite entanglement entropy
[…]
@@
Notice that this operation is different from evaluating the von Neumann entropy over the average state $|¯\rho_{t}=\overline{\ket{\psi(t)}\bra{\psi(t)}}$ which, besides that, would also not be a proper measure of the entanglement. Finally, we fix $\ell$ to be a fixed fraction of $L$ (in particular we consider either $\ell=L/2$ or $\ell=L/4$) and estimate the asymptotic long-time value of $\overline{S_{\ell}(t)}$ by performing a suitable time average:
@@
Here $t_{0},t_{f}]$ is an appropriate time window in which the behavior of $\overline{S_{\ell}(t)}$ has attained a steady-state value. We perform the average over quantum trajectories numerically, over a finite number of realizations that we fix as $N_{\rm r}=48$, when not differently specified. The error bars for our data are evaluated as the standard error (root-mean square deviation divided by $\sqrt{N_{\rm r}}$).
@@
$b\geq 1,$ $\qquad\text{for an area-law}.$
In what follows, we fit $\overline{S}_{\ell}$ versus $L$ for different fermionic models with Eq. (1). The most interesting result is that this function seems to fit our numerics very well, independently of the considered model. In particular, in the next section we specialize to integrable models, while in Sec. V we focus on nonintegrable models. For the latter, due to the small attainable system sizes ($L\lesssim 20$), it is more convenient to fit $\overline{S}_{\ell}$ versus the coupling $\gamma$ with the environment. We choose a generalized Lorentzian function and, in the end, we find that the parameters of the fit scale with the system size, in such a way that the form in Eq. (1) is recovered (details are provided in Sec. V.3). Finally, in Sec. VI, we switch to an free-fermion model with a more complicated geometry and a stroboscopic evolution, for which we consider the entanglement between two portions of a part of the whole system: as in that case the relevant part of the system is described by a mixed state, we resort to a proper entanglement monotone such as the FLN. To keep the presentation more accessible, we postpone all the required definitions to that section.